# On the Brun–Titchmarsh theorem

### v1

@@ 6.2. Transforming exponential sums
$\mathcal{R}(M,N)^{2}$ $\ll\frac{x^{2}}{MN^{2}}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ (n_{1}n_{2},q)=1\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}\sum_{\begin{subarray}{c}m\in\mathbf{Z}\\ (m,q)=1\end{subarray}}W\Big{(}\frac{m}{M}\Big{)}\mathrm{e}\Big{(}\frac{a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}m}}{q}\Big{)}.$
@@
$\varSigma^{*}$ $=\frac{M}{q}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1}\\ (n_{1}n_{2},q)=1\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}\sum_{1\leqslant|k|\leqslant K}\widehat{W}\Big{(}\frac{kM}{q}\Big{)}S(k,a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}};q).$
@@
    $\rho(u)$ $=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1},~{}(n_{1}n_{2},q)=1\\ (h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}}\equiv h\,(\operatorname{mod}{q})\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}.$
@@ 6.3. Bounding $\varSigma^{*}$ as bilinear forms
Note that the above congruence restriction $(h_{1}n_{2}-h_{2}n_{1})n_{1}^{\prime}n_{2}^{\prime}\equiv(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})n_{1}n_{2}\,(\operatorname{mod}{q})$ is equivalent to
(6.10) $(h_{1}n_{2}-h_{2}n_{1})n_{1}^{\prime}n_{2}^{\prime}=(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})n_{1}n_{2}+qt$
for some $t\in[-T,T]$ and $T:=1+4HN^{3}/q$.
Fix $n_{1}^{\prime},n_{2}^{\prime},h_{1},h_{2},t$ firstly, and choose a pair $(h_{1}^{*},h_{2}^{*})$ such that
(6.11) $(h_{1}n_{2}-h_{2}n_{1})n_{1}^{\prime}n_{2}^{\prime}=(h_{1}^{*}n_{2}^{\prime}-h_{2}^{*}n_{1}^{\prime})n_{1}n_{2}+qt,\ \ h_{1}^{*}n_{2}^{\prime}-h_{2}^{*}n_{1}^{\prime}\neq 0$
holds for some $n_{1},n_{2}.$ This may create $h_{1}^{\prime}=h_{1}^{*}+sn_{1}^{\prime}/(n_{1}^{\prime},n_{2}^{\prime})$ and $h_{2}^{\prime}=h_{2}^{*}+sn_{2}^{\prime}/(n_{1}^{\prime},n_{2}^{\prime})$ in (6.10) for some $w$ with $|s|\leqslant 2(n_{1}^{\prime},n_{2}^{\prime})H/N.$ Note that the number of such $s$ is at most $O(1+(n_{1}^{\prime},n_{2}^{\prime})H/N).$ Hence the problem reduces to counting solutions to the equation (6.11), which in fact implies
$(Cn_{1}-B)(Cn_{2}-A)=AB-Cqt,$
where
$A=-h_{1}n_{1}^{\prime}n_{2}^{\prime},\ \ B=h_{2}n_{1}^{\prime}n_{2}^{\prime},\ \ C=h_{1}^{*}n_{2}^{\prime}-h_{2}^{*}n_{1}^{\prime},\ \ ABC(AB-Cqt)\neq 0.$
Therefore, for any given $n_{1}^{\prime},n_{2}^{\prime},h_{1},h_{2},t$ therein and $h_{1}^{*},h_{2}^{*}$ as above, there exist at most $\tau(|AB-Cqt|)\ll(qHN)^{\varepsilon}$ choices of $n_{1},n_{2}.$ It then follows that
$\mathcal{B}_{2}$ $\ll H^{2}Tq^{\varepsilon}\mathop{\sum\sum}_{n_{1}^{\prime},n_{2}^{\prime}\sim N}(1+(n_{1}^{\prime},n_{2}^{\prime})H/N)$
$\ll(1+H/N)(1+HN^{3}/q)(HN)^{2}q^{\varepsilon}$
@@
(6.12) $M\leqslant q^{-1}x^{1-\varepsilon},\ \ N\leqslant\min\{x^{\frac{3}{4}}q^{-\frac{19}{16}},~{}x^{\frac{3}{8}}q^{-\frac{1}{2}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon}$
@@
(6.13) $M\leqslant q^{-1}x^{1-\varepsilon},\ \ N\leqslant\min\{x^{\frac{2}{3}}q^{-\frac{25}{24}},~{}x^{\frac{1}{3}}q^{-\frac{5}{12}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon}$
@@ 6.4. Bounding $\varSigma^{*}$ as trilinear forms
$\mathcal{T}_{2}(k_{1},k_{2})$ $\ll\frac{B}{q}(k_{1}-k_{2},q)+\frac{B}{\sqrt{q}}\frac{qC}{B}\Big{(}\frac{B}{C}+1\Big{)}^{\frac{1}{2}}q^{-\frac{1}{8}+\varepsilon}$
$\ll\frac{HN}{q}(k_{1}-k_{2},q)+\Big{(}\frac{H}{N}+1\Big{)}^{\frac{1}{2}}N^{2}q^{\frac{3}{8}+\varepsilon}.$
@@
$\mathcal{T}_{2}$ $\ll HN^{3}K+\frac{HNK^{2}}{q}+\Big{(}\frac{H}{N}+1\Big{)}^{\frac{1}{2}}(NK)^{2}q^{\frac{3}{8}+\varepsilon}$
(6.15) $\ll\frac{q^{2+\varepsilon}N^{4}}{x}+\frac{(qN)^{2+\varepsilon}}{xM}+\frac{N^{2}}{M^{2}}q^{\frac{19}{8}+\varepsilon}$
@@
$\varSigma^{*}$ $\ll\Big{(}\frac{q^{\frac{3}{2}}M^{2}N^{4}}{x^{\frac{3}{2}}}+\frac{(qM)^{\frac{3}{2}}N^{4}}{x^{\frac{3}{2}}}+\frac{q^{\frac{27}{16}}MN^{3}}{x}\Big{)}q^{\varepsilon}.$
@@
$\mathcal{R}(M,N)^{2}$ $\ll(xqM+x^{\frac{1}{2}}q^{\frac{3}{2}}MN^{2}+x^{\frac{1}{2}}q^{\frac{3}{2}}M^{\frac{1}{2}}N^{2}+xq^{\frac{27}{16}}N+qM^{2}N^{2})q^{\varepsilon}.$
@@
(6.16) $M\leqslant q^{-1}x^{1-\varepsilon},\ \ N\leqslant\min\{xq^{-\frac{27}{16}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon}$
@@
In view of (6.12), (6.13) and (6.16), we find (6.1) is valid if taking $M=q^{-1}x^{1-\varepsilon}$ and
$N=\begin{cases}\min\{xq^{-\frac{27}{16}},~{}(x/q)^{\frac{1}{4}}\}\}x^{-\varepsilon},\ \ &\varpi\in[1/2,~{}32/61[,\\ \min\{x^{\frac{2}{3}}q^{-\frac{25}{24}},~{}x^{\frac{1}{3}}q^{-\frac{5}{12}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon},&\varpi\in[32/61,~{}7/13[,\\ \min\{x^{\frac{3}{4}}q^{-\frac{19}{16}},~{}x^{\frac{3}{8}}q^{-\frac{1}{2}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon},&\varpi\in[7/13,~{}4/7[.\end{cases}$
@@
with $\varepsilon\gg\eta.$ In particular, for $\varpi\geqslant\frac{1+\kappa-\lambda}{2+2\kappa-\lambda}$, we may take
@@
7. Proof of Theorem 3.1
@@ 7.1. Preparations
Denote by $\omega(n)$ the number of distinct prime factors of $n$. We first recall a very crucial observation due to Fouvry [Fo84, Lemme 6].
@@
$-β\displaystyle\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ $\ll\|\boldsymbol{\alpha}\|(HNS)^{\frac{1}{2}}((HNS)^{\frac{1}{4}}+(NS)^{\frac{1}{2}})(\log(NS))^{2\xi}\Bigg{[}NS(MR)^{\frac{1}{2}}\frac{(|a|+\frac{MNRS}{H})^{\vartheta}}{1+(\frac{|a|H}{MNRS})^{2}}$
@@
$\ \ \ \ +\|\boldsymbol{\alpha}\|(HMNRS)^{\frac{1}{2}+\varepsilon}(\log(NS))^{2\xi},$
where the implied constant depends only on $\varepsilon,$ and $\vartheta\leqslant 7/64$ is the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$
@@
According to Lemma 7.1, we split the sums over $n,s$ into at most $O((\log NS)^{2\xi})$ sub-sums of the shape
@@
$\ll\|\boldsymbol{\alpha}\|^{2}HMNRS(HNS)^{\varepsilon}.$
Denote by $\mathcal{Q}^{*}$ the contribution from off-diagonal terms, so that
(7.3) $|\mathcal{Q}^{*}|^{2}$ $\leqslant\mathcal{Q}_{1}+O(\|\boldsymbol{\alpha}\|^{2}HMNRS(HNS)^{\varepsilon}).$
@@
For given $n_{1},n_{2},s_{1},s_{2}$, we introduce a coefficient $\boldsymbol{\gamma}_{n_{1},n_{2},s_{1},s_{2}}=(\gamma_{k;n_{1},n_{2},s_{1},s_{2}})$, indexed by $k$, as
$\gamma_{k;n_{1},n_{2},s_{1},s_{2}}$ $=\mathop{\sum\sum\sum\sum\sum\sum}_{\begin{subarray}{c}h_{1},h_{2}\leqslant H,~{}n_{1},n_{2}\leqslant N,~{}s_{1},s_{2}\leqslant S\\ (n_{1}n_{2},s_{1}s_{2})=1,~{}(n_{1},s_{1}),(n_{2},s_{2})\in\mathscr{F}_{\xi}^{*}(N,S)\\ h_{1}n_{2}s_{2}-h_{2}n_{1}s_{1}=k\end{subarray}}\beta_{h_{1},n_{1},s_{1}}\overline{\beta_{h_{2},n_{2},s_{2}}}.$
@@
$(q,r,s)\leftarrow(a,n_{1}n_{2},s_{1}s_{2}),\ \ (c,m,n)\leftarrow(r,m,k),\ \ (C,M,N)\leftarrow(R,M,HNS),$
@@
$=\sqrt{MN^{2}RS^{2}}\frac{(|a|+\frac{MNRS}{H})^{\vartheta}}{1+(\frac{|a|H}{MNRS})^{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{R}{MN^{2}}\Big{)}^{\frac{1}{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big{)}^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}.$
@@
×$\ \ \ \ \times\Big{(}1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big{)}^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}\Bigg{]},$

### v3

@@ 6.2. Transforming exponential sums
$\mathcal{R}(M,N)^{2}$ $\ll\frac{x^{2}}{MN^{2}}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ (n_{1}n_{2},q)=1\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}\sum_{\begin{subarray}{c}m\in\mathbf{Z}\\ (m,q)=1\end{subarray}}$
×$\ \ \ \ \times W\Big{(}\frac{m}{M}\Big{)}\mathrm{e}\Big{(}\frac{a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}m}}{q}\Big{)}.$
@@
$\varSigma^{*}$ $=\frac{M}{q}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1}\\ (n_{1}n_{2},q)=1\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}\sum_{1\leqslant|k|\leqslant K}\widehat{W}\Big{(}\frac{kM}{q}\Big{)}$
×$\ \ \ \ \times S(k,a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}};q).$
@@
    $\rho(u)$ $=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1},~{}(n_{1}n_{2},q)=1\\ (h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}}\equiv u\,(\operatorname{mod}{q})\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}.$
@@ 6.3. Bounding $\varSigma^{*}$ as bilinear forms
Note that the above congruence restriction $(h_{1}n_{2}-h_{2}n_{1})n_{1}^{\prime}n_{2}^{\prime}\equiv(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})n_{1}n_{2}\,(\operatorname{mod}{q})$ implies
$(h_{1}n_{2}-h_{2}n_{1})n_{1}^{\prime}n_{2}^{\prime}=(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})n_{1}n_{2}+qt,\ \ t\in\mathbf{Z}\cap[-T,T]$
with $T:=4HN^{3}/q.$ Denote by $\mathcal{B}_{2,0}$ and $\mathcal{B}_{2,1}$ the contributions from $t=0$ and $t\neq 0$, respectively.
We now consider $\mathcal{B}_{2,1}$, which vanishes unless $T\geqslant 1$. Picking out the g.c.d. of $n_{1},n_{2}$ and that of $n_{1}^{\prime},n_{2}^{\prime}$, each of which divides $t\neq 0$, we obtain
$\mathcal{B}_{2,1}$ $\ll\log^{2}N\sup_{D,D^{\prime}\leqslant 2N}\mathcal{B}_{2,1}(D,D^{\prime})$
with
$\mathcal{B}_{2,1}(D,D^{\prime})$ $=\mathop{\sum\sum\sum\sum\sum\sum\sum\sum\sum\sum\sum}_{\begin{subarray}{c}d\sim D,~{}d^{\prime}\sim D^{\prime},~{}n_{1},n_{2}\sim N/d,~{}n_{1}^{\prime},n_{2}^{\prime}\sim N/d^{\prime},~{}1\leqslant|h_{1}|,|h_{1}^{\prime}|,|h_{2}|,|h_{2}^{\prime}|\leqslant H,~{}1\leqslant|t|\leqslant T/DD^{\prime}\\ (n_{1},n_{2})=(n_{1}^{\prime},n_{2}^{\prime})=1\\ (h_{1}n_{2}-h_{2}n_{1})(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})\neq 0\end{subarray}}1,$
where the summations are also restricted to
$(h_{1}n_{2}-h_{2}n_{1})d^{\prime}n_{1}^{\prime}n_{2}^{\prime}=(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})dn_{1}n_{2}+qt.$
Given an integer $\ell$ with $0<|\ell|\leqslant 2HN/D,$ consider the equations
(6.10) $h_{1}n_{2}-h_{2}n_{1}=\ell,$
and
(6.11) $\ell d^{\prime}n_{1}^{\prime}n_{2}^{\prime}=(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})dn_{1}n_{2}+qt.$
If fixing $n_{1},n_{2}$ with $(n_{1},n_{2})=1$, the number of pairs $(h_{1},h_{2})$ with $1\leqslant|h_{1}|,|h_{2}|\leqslant H$ satisfying (6.10) is $\ll 1+H/n_{1}\ll 1+DH/N.$ Regarding the equation (6.11), we turn to consider the congruence equation
(6.12) $(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})dn_{1}n_{2}\equiv-qt\,(\operatorname{mod}{d^{\prime}n_{1}^{\prime}n_{2}^{\prime}}).$
Given $h_{1}^{\prime},h_{2}^{\prime},n_{1}^{\prime},n_{2}^{\prime},t$, the number of tuples $(d,n_{1},n_{2})$ satisfying (6.12) with $d\sim D,n_{1},n_{2}\sim N/d$ is
$\ll\Big{(}\frac{N^{2}}{D}\frac{(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime},d^{\prime}n_{1}^{\prime}n_{2}^{\prime})}{d^{\prime}n_{1}^{\prime}n_{2}^{\prime}}+1\Big{)}N^{\varepsilon}\ll\Big{(}\frac{D^{\prime}}{D}(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime},d^{\prime}n_{1}^{\prime}n_{2}^{\prime})+1\Big{)}N^{\varepsilon},$
where the factor $N^{\varepsilon}$ comes from the bound $\tau_{3}(m)\ll m^{\varepsilon}.$
Collecting all above arguments, we find
$\mathcal{B}_{2,1}(D,D^{\prime})$ $\ll q^{\varepsilon}(1+DH/N)\frac{T}{DD^{\prime}}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}d^{\prime}\sim D^{\prime},~{}n_{1}^{\prime},n_{2}^{\prime}\sim N/d^{\prime},~{}1\leqslant|h_{1}^{\prime}|,|h_{2}^{\prime}|\leqslant H\\ h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime}\neq 0,~{}(n_{1}^{\prime},n_{2}^{\prime})=1\end{subarray}}\Big{(}\frac{D^{\prime}}{D}(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime},d^{\prime}n_{1}^{\prime}n_{2}^{\prime})+1\Big{)}$
$\ll q^{\varepsilon}(1+DH/N)\frac{T}{DD^{\prime}}(HN)^{2},$
so that
$\mathcal{B}_{2,1}$ $\ll(1+H/N)(HN)^{2}Tq^{\varepsilon}.$
Following similar arguments, we obtain
$\mathcal{B}_{2,0}$ $\ll(1+H/N)(HN)^{2}q^{\varepsilon}.$
Therefore, we conclude
$\mathcal{B}_{2}$ $\ll(1+H/N)(1+T)(HN)^{2}q^{\varepsilon}$
@@
(6.13) $M\leqslant x^{1-\varepsilon}q^{-1},\ \ N\leqslant\min\{x^{\frac{3}{4}}q^{-\frac{19}{16}},~{}x^{\frac{3}{8}}q^{-\frac{1}{2}},~{}q^{\frac{1}{4}}\}x^{-\varepsilon}$
@@
(6.14) $M\leqslant x^{1-\varepsilon}q^{-1},\ \ N\leqslant\min\{x^{\frac{2}{3}}q^{-\frac{25}{24}},~{}x^{\frac{1}{3}}q^{-\frac{5}{12}},~{}q^{\frac{1}{4}}\}x^{-\varepsilon}$
@@ 6.4. Bounding $\varSigma^{*}$ as trilinear forms
$\mathcal{T}_{2}(k_{1},k_{2})$ $\ll\frac{BC}{q}(k_{1}-k_{2},q)+\frac{B}{\sqrt{q}}\frac{qC}{B}\Big{(}\frac{B}{C}+1\Big{)}^{\frac{1}{2}}q^{-\frac{1}{8}+\varepsilon}$
$\ll\frac{HN^{3}}{q}(k_{1}-k_{2},q)+\Big{(}\frac{H}{N}+1\Big{)}^{\frac{1}{2}}N^{2}q^{\frac{3}{8}+\varepsilon}.$
@@
$\mathcal{T}_{2}$ $\ll HN^{3}K+HN^{3}K^{2}q^{-1+\varepsilon}+\Big{(}\frac{H}{N}+1\Big{)}^{\frac{1}{2}}(NK)^{2}q^{\frac{3}{8}+\varepsilon}$
(6.16) $\ll\frac{N^{4}}{x}q^{2+\varepsilon}+\frac{N^{2}}{M^{2}}q^{\frac{19}{8}+\varepsilon}$
@@
$\varSigma^{*}$ $\ll\Big{(}\frac{q^{\frac{3}{2}}M^{2}N^{4}}{x^{\frac{3}{2}}}+\frac{q^{\frac{27}{16}}MN^{3}}{x}\Big{)}q^{\varepsilon}.$
@@
$\mathcal{R}(M,N)^{2}$ $\ll(xqM+x^{\frac{1}{2}}q^{\frac{3}{2}}MN^{2}+xq^{\frac{27}{16}}N+qM^{2}N^{2})q^{\varepsilon}.$
@@
(6.17) $M\leqslant x^{1-\varepsilon}q^{-1},\ \ N\leqslant\min\{xq^{-\frac{27}{16}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon}$
@@
In view of (6.13), (6.14) and (6.17), we find (6.1) is valid if taking $M=x^{1-\varepsilon}q^{-1}$ and
$N=\begin{cases}\min\{xq^{-\frac{27}{16}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon},\ \ &\varpi\in[1/2,~{}32/61[,\\ \min\{x^{\frac{2}{3}}q^{-\frac{25}{24}},~{}x^{\frac{1}{3}}q^{-\frac{5}{12}}\}x^{-\varepsilon},&\varpi\in[32/61,~{}7/13[,\\ \min\{x^{\frac{3}{4}}q^{-\frac{19}{16}},~{}x^{\frac{3}{8}}q^{-\frac{1}{2}}\}x^{-\varepsilon},&\varpi\in[7/13,~{}4/7[.\end{cases}$
@@
Remark 5.
Each of the above choices of $N$ fails to improve the work of Iwaniec [Iw82] when $\varpi=1/2.$ This is due to the fact that we have explored the bilinear or trilinear structures in $\varSigma^{*},$ which requires that $K>q^{4\varepsilon}.$ This is impossible by recalling that $K=M^{-1}q^{1+\varepsilon}$ and $M=x^{1-\varepsilon}q^{-1}.$ We will beat this barrier in a forthcoming work ([XZ25]), i.e., we are able to improve Iwaniec [Iw82] when $\varpi=1/2.$
@@
with $\varepsilon\gg\eta.$ In particular, for $\varpi\geqslant\frac{1+\kappa-\lambda}{2+2\kappa-\lambda}$, the choice $N=x^{\frac{1+\kappa-\lambda}{2}-\varepsilon}q^{-\frac{1+2\kappa-\lambda}{2}}$ is admissible, so that we may take
@@
7. Proof of Theorems 2.1 and 2.2
The proofs of Theorems 2.1 and 2.2 would follow the same line, except that Lemma 2.3 applies in the proof of Theorem 2.1, and we utilize convolution structures of the coefficients when proving Theorem 2.2, so that Kloostermania (Lemma 2.4) applies.
@@ 7.1. Preparations
Denote by $\omega(n)$ the number of distinct prime factors of $n$. We first recall a very crucial observation due to Fouvry [Fo84a, Lemme 6].
@@
Let $H,M,N,R,S,\xi\geqslant 1$ and $a\in\mathbf{Z}^{*}.$ We assume that $\beta_{h,n,s}=0$ unless $(n,s)\in\mathscr{F}_{\xi}(N,S)$ and $\mu(ns)\neq 0.$ Then we have
@@
[…]
where the implied constant depends only on $\varepsilon,\xi,$ and $\vartheta\leqslant 7/64$ is the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$
@@
[…]
@@
$\ll\|\boldsymbol{\alpha}\|^{2}(HMNRS)^{1+\varepsilon}.$
Denote by $\mathcal{Q}_{1}$ the contribution from off-diagonal terms, so that
(7.3) $|\mathcal{Q}^{*}|^{2}$ $\leqslant\mathcal{Q}_{1}+O(\|\boldsymbol{\alpha}\|^{2}(HMNRS)^{1+\varepsilon}).$
@@
For given $n_{1},n_{2},s_{1},s_{2}$ with $(n_{1}n_{2},s_{1}s_{2})=1$ and $(n_{1},s_{1}),(n_{2},s_{2})\in\mathscr{F}_{\xi}^{*}(N,S)$, we introduce a coefficient $\boldsymbol{\gamma}_{n_{1},n_{2},s_{1},s_{2}}=(\gamma_{k;n_{1},n_{2},s_{1},s_{2}})$, indexed by $k$, as
$\gamma_{k;n_{1},n_{2},s_{1},s_{2}}$ $=\mathop{\sum\sum}_{\begin{subarray}{c}h_{1},h_{2}\leqslant H\\ h_{1}n_{2}s_{2}-h_{2}n_{1}s_{1}=k\end{subarray}}\beta_{h_{1},n_{1},s_{1}}\overline{\beta_{h_{2},n_{2},s_{2}}}.$
@@
$(a,r,s)\leftarrow(a,n_{1}n_{2},s_{1}s_{2}),\ \ (c,m,n)\leftarrow(r,m,k),\ \ (C,M,N)\leftarrow(R,M,HNS),$
@@
$=\sqrt{MN^{2}RS^{2}}\frac{(|a|+\frac{MNRS}{H})^{\vartheta}}{1+(\frac{|a|H}{MNRS})^{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{R}{MN^{2}}\Big{)}^{\frac{1}{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big{)}^{\frac{1}{2}}$
$\ \ \ \ +S^{-2}M(HNS)^{\frac{1}{2}}.$
@@
×$\ \ \ \ \times\Big{(}1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big{)}^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}\Bigg{]}(HMNRS)^{\varepsilon},$